# MultiBLiMP 1.0: A Massively Multilingual Benchmark of Linguistic Minimal Pairs

### v1

@@ Acknowledgments
We would like to thank Amir Kargaran for providing the language frequencies in CommonCrawl. We further thank Ahmet Üstün, Charlotte Pouw, Jirui Qi, Kanishka Misra and Will Merrill, as well as the GroNLP group for helpful feedback. Jaap Jumelet and Arianna Bisazza were supported by NWO grant VI.Vidi.221C.009. Leonie Weissweiler was supported by a postdoctoral fellowship from the German Research Foundation (DFG). This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-13403.
@@ A.1 Excluded Treebanks and Criteria
We exclude a number of treebanks for various reasons. For low-resource languages, treebanks are only excluded where absolutely necessary, for high-resource languages, treebanks may be excluded to make the resulting genres and annotation schemata more uniform. An overview of the treebanks we removed is provided in Table 4. We remove diacritics from Latin, Slovenian, and Western Farsi, and cantillations from Biblical Hebrew to ensure consistency between their UD and UM data sources.. We transliterate Uyghur from Arabic to Latin, Sanskrit from Latin to Devanagari, and Tatar from Latin to Cyrillic using the official guidelines https://suzlek.antat.ru/about/TAAT2019/8.pdf. We skip all UD languages that do not have an ISO-639-3 code: code-switched Frisian Dutch, Turkish German, and Telugu English, as well as Cappadocian, Maghrebi Arabic French, Pomak.
@@ B.1 Language Distribution
B.2 Best Models per Language

### v2

@@ 8 Discussion and Conclusion
Limitations
Our minimal pair creation procedure is limited by the size, diversity, and annotation quality of the resources which we rely on, namely UD and UniMorph. Extending our procedure to more complex phenomena may also be challenging, since it requires one to define grammaticality violations in terms of morphological inflections. We are further limited by the sheer number of languages in our benchmark, which makes manual evaluation cost-prohibitive, except for spot checks for a subset of languages. We see MultiBLiMP as a continuous effort, for which we have already taken feedback from language experts in our network and adjusted individual languages accordingly.
@@ Acknowledgments
We would like to thank Amir Kargaran for providing the language frequencies in CommonCrawl. We further thank Ahmet Üstün, Charlotte Pouw, Jirui Qi, Omer Goldman, Kanishka Misra and Will Merrill, as well as the GroNLP group for helpful feedback. Jaap Jumelet and Arianna Bisazza were supported by NWO grant VI.Vidi.221C.009. Leonie Weissweiler was supported by a postdoctoral fellowship from the German Research Foundation (DFG). This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-13403.
@@ A.1 Excluded Treebanks and Criteria
We exclude a number of treebanks for various reasons. For low-resource languages, treebanks are only excluded where absolutely necessary, for high-resource languages, treebanks may be excluded to make the resulting genres and annotation schemata more uniform. An overview of the treebanks we removed is provided in Table 4. We remove diacritics from Latin, Slovenian, and Western Farsi, and cantillations from Biblical Hebrew to ensure consistency between their UD and UM data sources. We transliterate Uyghur from Arabic to Latin, Sanskrit from Latin to Devanagari, and Tatar from Latin to Cyrillic using the official guidelines https://suzlek.antat.ru/about/TAAT2019/8.pdf. We skip all UD languages that do not have an ISO-639-3 code: code-switched Frisian Dutch, Turkish German, and Telugu English, as well as Cappadocian, Maghrebi Arabic French, Pomak.
@@ B.1 Language Distribution
B.2 Significantly Best Models per Language